# SA-Person: Text-Based Person Retrieval with Scene-aware Re-ranking

### v2

@@ 4.4 Ablation Studies
We conduct ablation studies to evaluate the effectiveness of the proposed SceneRanker, focusing on (1) the prompt design strategy, (2) the choice of multimodal large language models (MLLMs), and (3) the candidate size in the first-stage retrieval.
@@
Table 4: Comparison the impact of using different MLLMs into the SceneRanker. Due to varying MLLM capabilities in processing structured prompts, we adopt Naive Prompt across all models to ensure fair comparison.
@@
IRRA (ViT-L/14) 63.37 84.54 88.67 72.71
QwenVL2.5 bai2025qwen2 3 Billion 54.57 78.23 88.67 65.04
8 Billion 75.07 86.34 88.67 80.45
MiniCPM-V-2_6 yao2024minicpm 8 Billion 71.33 84.66 88.67 77.78
InternVL2.5 chen2024expanding 1 Billion 3.01 80.14 88.67 26.71
@@
Prompt Design R@1 R@5 R@10 mAP 63.37 84.54 88.67 72.71 Naive Prompt 77.55 87.42 88.67 82.33 Box Overlay Prompt 76.69 87.29 88.67 81.85 Box Embedded Prompt 78.34 87.54 88.67 82.86
Table 5: Impact of visual prompt design on the performance of SceneRanker. IRRA(ViT-L/14) is adopted as the baseline for comparison, without prompt design.
1) Naive Prompt (NP), which provides only the textual description and the full-scene image without explicit region information; 2) Box Overlay Prompt (BOP), which visually highlights the candidate region with a red rectangle on the full-scene image; 3) Box Embedded Prompt (BEP), which includes the bounding box coordinates of the candidate region in the multimodal prompt using <box> tags. The complete designs of the three prompt strategies are provided in Appendix A.5. As shown in Table 5, incorporating box embedded prompt into the proposed SceneRanker outperforms naive prompt by 0.79% at R@1 and 0.53% at mAP, confirming that explicitly encoding region locations strengthens the model’s spatial grounding. Compared to the box overlay prompt, box embedded prompt achieves an even greater improvement of 1.65% at R@1 and 1.01% at mAP, suggesting that structured coordinate inputs are more effective than visual overlays. Notably, BOP even play worse than NP by 0.86% at R@1 and 0.48% at mAP, indicating that direct visual highlighting may introduce visual distractions that interfere with the model’s scene understanding.
Different Multimodal Large Language Models. We evaluate the effect of integrating different 8B-scale MLLMs into the proposed SceneRanker, using the same naive prompt for fair comparison. As shown in Table 4, all three models, Qwen-VL 2.5, MiniCPM-V, and InternVL 2.5, consistently improve over the IRRA baseline, with respective R@1 gains of 11.7%, +7.96%, and +14.18%. In addition, we observe a strong correlation between model capacity and performance. Taking InternVL 2.5 as an example, scaling from 4B to 8B brings further improvement, with R@1 increasing from 75.21% to 77.55%. In contrast, smaller models below 4B such as InternVL 1B and 2B show significantly lower performance, indicating insufficient multimodal understanding. These findings suggest that various MLLMs are well-suited for SceneRanker, and its performance benefits from stronger vision-language reasoning capabilities. Importantly, this holds true regardless of the specific architecture, provided the model has sufficient capacity.
Different Candidate Sizes. In the proposed two-stage framework SA-Person, the candidate set produced by the appearance-based retrieval stage serves as the input for the SceneRanker to perform scene-aware reranking. The size of this set ($K$) directly affects the trade-off between including the correct match and introducing noisy distractors. To better understand this trade-off and determine an optimal value for ($K$), we conduct a parameter analysis to systematically
@@
investigate the impact of different candidate set sizes on the final retrieval performance. As shown in Figure  3, the performance of SceneRanker exhibits a trend of first increasing and then decreasing as the candidate size changes. The best performance is achieved when the candidate size is set to 10. Smaller candidate sets are more likely to exclude the ground-truth instance, while excessively large candidate sets introduce more irrelevant or noisy candidates. Both scenarios negatively impact performance. These results indicate that a moderate candidate pool is optimal, as it balances retaining the ground-truth instance while minimizing noise, ultimately enhancing retrieval accuracy.
4.5 Qualitative Results
To further validate the effectiveness of SA-Person, we present a qualitative comparison in Figure 4. IRRA focuses solely on pedestrian appearance, retrieving candidates with similar coarse attributes, but fails to utilize scene information and thus struggles to distinguish individuals in different contexts. LongCLIP, on the other hand, leverages scene-level information to retrieve images consistent with the described environment, but lacks fine-grained attention to pedestrian appearance and often returns individuals who differ significantly from the target identity. In comparison, SA-Person first retrieves candidates based on pedestrian appearance and then employs an MLLM to understand scene context and re-rank these candidates accordingly. This approach enables SA-Person to more accurately align with both the appearance and context described in the query, leading to significantly improved retrieval accuracy over existing baselines.
Figure 4: Qualitative comparison of retrieval results across three methods. The leftmost column shows the textual query and its corresponding ground-truth image. The top-6 retrieval results are presented from top to bottom for IRRA (ViT-L/14), LongCLIP (ViT-L/14), and SA-Person (ours).
5 Conclusion
In this paper, we tackle the challenging task of person retrieval by leveraging multimodal information from both pedestrian appearance and global scene context. Specifically, we make two key contributions. First, we present ScenePerson-13W, the first large-scale dataset to offer comprehensive multimodal annotations, integrating fine-grained pedestrian appearance descriptions with rich scene-level contextual information within full-scene images. Second, we propose SceneRanker, a novel scene-aware re-ranking method that exploits multimodal cues from the entire scene to effectively refine retrieval results. Building upon these contributions, we develop SA-Person, a two-stage framework that initially grounds discriminative appearance features in cropped pedestrian regions and subsequently applies SceneRanker, a training-free multimodal large language model—to perform scene-aware ranking based on holistic image context. Extensive experimental results demonstrate that SA-Person significantly outperforms existing baselines, underscoring the critical importance of integrating both local appearance cues and global scene context for robust and effective text-based person retrieval in complex real-world scenarios.
Limitations. While our work advances scene-aware person retrieval, it has two main limitations. First, the query descriptions in ScenePerson-13W are generated by MLLM, which may slightly deviate from natural human expression in retrieval scenarios. Second, our SceneRanker builds on off-the-shelf MLLMs without task-specific training, leaving room for improvement through fine-tuning to enhance alignment with retrieval intent and strengthen scene context understanding.
Appendix A Appendix of SA-Person: Text-Based Person Retrieval with Scene-aware Re-ranking
In the appendix, we provide the following materials:
  • Section A.1 provides a discussion of related work on text-based person retrieval and multimodal large language models.
  • Section A.2 analyzes the broader impact of scene-aware person retrieval, addressing societal benefits and ethical considerations.
  • Section A.3 details the construction and description generation processes for the SCENEPERSON-13W dataset.
  • Section A.4 presents supplemental results based on varisou retrieval methods and prompt designs, and visualizations of SceneRanker’s performance in Section 4.4 and 4.5 in the main paper.
  • Section A.5 describes prompt designs and instructional templates for SceneRanker in Section 4.4 in the main paper.
A.1 Related works
Text-Based Person Retrieval. Text-Based Person Retrieval (TBPR) aims to retrieve specific individuals based on natural language descriptions, typically from a gallery of cropped person images [45, 19, 55, 26, 25]. These methods focus on fine-grained cross-modal alignment between visual appearances and textual attributes, such as clothing, gender, or accessories [34, 55, 37, 38, 51]. With the advent of vision-language models like CLIP [31], TBPR performance has significantly improved. For instance, IRRA [15] introduces a multi-modal interaction encoder for implicit relational reasoning, and TBPS-CLIP [5] proposes a simple yet strong baseline through enhanced training strategies. To address noise in cross-modal data, RaSa [3] proposes sensitivity-aware and relation-aware representation learning for robustness against weakly matched pairs; MARS [10] incorporates visual reconstruction and attribute-sensitive losses to mitigate inter-identity noise and intra-identity variations; RDE [30] introduces dual embeddings to handle under-correlated or mislabeled image-text pairs under noisy conditions. Beyond cropped-image retrieval, some methods [50, 35, 42] retrieve individuals directly from unprocessed full-scene images, where individuals need to be located and identified in complex visual environments. Specifically, Zhang et al. [50] integrates textual cues into region proposals, enabling text descriptions to guide the detection process. Su et al. [35] introduces a memory-aided coarse-to-fine alignment strategy, using memory banks to enhance cross-modal contrastive learning. Yan et al. [42] explores multimodal fusion by combining textual and visual features with precise segmentation techniques to boost retrieval performance. While these methods often overlook the rich contextual information present in the surrounding scene, which could offer valuable clues for more accurate and robust person retrieval. Our SA-Person framework distinguishes itself by explicitly integrating scene context with appearance-based retrieval, leveraging textual descriptions of both individuals and their surroundings to enhance accuracy and disambiguate similar-looking candidates.
Multimodal Large Language Models for Scene and Region-Level Reasoning. Recent advances in Multimodal Large Language Models (MLLMs) enabled powerful visual reasoning across complex scenes. Current efforts primarily progresses along two directions: comprehensive scene understanding using global image-text inputs [7, 8, 2, 36, 28, 11], and fine-grained region-level grounding [13, 54, 29, 6, 27, 32]. For comprehensive scene understanding, recent works [28, 11] enable MLLMs to explicitly represent structured and compositional semantics. For instance, Mitra et al. [28] proposes generating scene graphs via MLLMs and incorporating them into prompts, facilitating compositional reasoning without relying on annotated scene graphs; MLLM-SUL [11] introduces a dual-branch encoder and a fine-tuned LLaMA [36] model to jointly perform scene description in traffic scenarios. To support region-level grounding, recent studies [13, 54, 6] enhance MLLMs with explicit spatial conditioning. For example, RegionGPT [13] and RegionBLIP [54] introduce region-wise feature pooling within bounding boxes or masks to extract localized representations for improved spatial reasoning; KOSMOS-2 [29] and Shikra [6] propose to encode spatial positions as discrete tokens or numeric coordinates to support referential understanding; GLaMM [32] facilitates grounded interaction through joint generation of text and segmentation masks. In this work, we explore the application of MLLMs to fine-grained person retrieval in complex scenes, where both textual descriptions and full-scene images are jointly leveraged to assess and rank visual relevance.
A.2 Broader impacts
The SA-Person framework and SCENEPERSON-13W dataset advance text-based person retrieval, improving public safety, smart city accessibility, and multimodal AI research. However, their surveillance applications risk privacy violations and biases from YouTube-derived data. To mitigate, restricting access with ethical guidelines ensures only responsible, research-focused use, preventing misuse by malicious actors and aligning with legal standards. Enhancing robustness through fine-tuning improves model accuracy across diverse scenarios, reducing errors and increasing reliability, thus maximizing societal benefits while ensuring ethical deployment.
Figure 5: Overview of person retrieval limitation and scene-aware insights. (a) Illustration of appearance-based retrieval limitation, where multiple candidates match the query “A girl wearing a green and white striped shirt is smiling” due to similar appearances. (b) Illustration of scene-aware reranking, ranking candidate images by aligning scene context like “located in a restaurant, near a window with red text” within the query. (c) Wordcloud of contextual elements in ScenePerson-13W, highlighting diverse scene descriptors. (d) Examples from the ScenePerson-13W dataset, showcasing detailed scene and relational descriptions for person retrieval.
Figure 6: Example instances from the ScenePerson-13W dataset. Descriptions in green text detail pedestrian appearance, while those in blue text describe their interactions within complex scenes. To clearly identify the target pedestrian associated with each textual description in scenes containing multiple pedestrians, the target individual is highlighted with a red bounding box in these visualizations.
A.3 More Details of ScenePerson-13W
The ScenePerson-13W dataset addresses critical challenges in scene-aware person retrieval by integrating detailed pedestrian appearance with rich scene context, overcoming limitations of appearance-only methods, as illustrated in Figure 5. This section provides additional details on the construction pipeline of the SCENEPERSON-13W dataset, with a focus on specific methodologies and confidence thresholds to ensure reproducibility and clarity.
Pedestrian Detection and Tracking. Using the YOLOv11x, we detect and track pedestrians in the QVHighlights video dataset [17] with a confidence threshold of 0.6 to balance precision and recall.
Completeness Filtering. To ensure quality, incomplete detections are filtered using MediaPipe’s pose estimation [1] with a minimum detection confidence of 0.65. Head detection requires at least one keypoint (e.g., nose, eyes) with visibility over 0.4, and shoulders need visibility over 0.6 for at least one shoulder. Only images with a head and one shoulder are retained.
Deduplication. Redundancy is reduced using CLIP [31] to compute cosine similarity, normalized to the range [0, 1], for cropped pedestrian regions and full-scene images. Images with similarity score exceeding 0.91 are deemed redundant, and one instance is kept, yielding 131,438 unique pedestrian instances for diversity.
Testing Set Refinement. To create a fair and robust testing set of 10,000 images, we address redundancy from frame-based tracking, where pedestrians appear across consecutive frames with minor variations. We apply CLIP-based visual clustering using the clip-vit-base-patch32 [31] model, with a similarity threshold of 0.88. Images are clustered via agglomerative clustering with average linkage, and one representative image per cluster is selected. This deduplicated testing set balances diversity and manageability, preserving complex scene and identity distributions for meaningful retrieval evaluation.
Description Generation. Textual descriptions for each of the 131,438 pedestrian instances are generated using InternVL2.5-8B [8]. The model is prompted with a full-scene image, highlighting the target pedestrian with a red bounding box, using the following prompt:
Generate a description of the individual within the red bounding box and the connection with the surroundings, in a style similar to the template: template. Ensure that the description is concise within 50 words and accurate. Do not include text details or labels, and strictly avoid any punctuation other than commas and periods.
The template guides the model to describe: (1) the individual’s appearance, (2) their location in the scene, and (3) relationships with surroundings. Specifically, the template is defined as:
A [man/woman/girl/boy] wearing [describe basic clothing, notable accessories, and physical attributes] is [describe their action or posture]. This person is located [describe their location, highlighting any landmarks, structures, or recognizable surroundings]. Next to/to the left of/to the right of/behind/in front of/under/above/around [the man/woman/etc.], there is/are [briefly describe nearby people, objects, or structures, such as another person in red holding something, a group of people talking, or a pedestrian bridge].
This structured approach ensures descriptions capture both pedestrian details and contextual cues, as illustrated in Figure 6.
Dataset Examples. More examples of ScenePerson-13W are shown in Figure 6. showcasing the dataset’s rich scene and relational descriptions. These examples highlight the dataset’s ability to provide fine-grained appearance details and diverse scene contexts, addressing challenges like those in Figure 5, where scene-aware insights disambiguate similar-looking individuals.
A.4 Experiments and Visualisation
In this section, we provide deeper insights into our experimental results, ablation findings, and qualitative visualizations.
A.4.1 SceneRanker Performance Across Retrieval Methods and Prompt Designs
Table 6: Performance of SceneRanker with various text-based person retrieval methods under different prompt designs. Red numbers indicate the improvement brought by each prompt design.

### v3

@@ VI-D Ablation Studies
We conduct ablation studies to evaluate the effectiveness of the proposed SceneRanker, focusing on (1) the prompt design strategy, (2) the choice of MLLMs, and (3) the candidate size in the first-stage retrieval.
TABLE V: Performance of SceneRanker with various text-based person retrieval methods under different prompt designs. Red numbers indicate the improvement brought by each prompt design.
Methods R@1 R@5 R@10 mAP
TBPS-CLIP [cao2024empirical] 50.29 75.24 81.58 61.54
IRRA(ViT-B/16) [jiang2023cross] 60.57 83.35 87.86 70.65
IRRA(ViT-L/14) [jiang2023cross] 63.37 84.54 88.67 72.71
RDE [qin2024noisy] 61.26 83.50 88.23 71.17
+NP $77.13_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.87}}$ $6.83_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.33}} 88$$8.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.87_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.70}}$
+BOP $76.29_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.03}}$ $\textbf{86.88}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.38}}$ $88.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.46_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.29}}$
+BEP $\textbf{77.58}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.32}}$ ${86.84}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.34}}$ $\textbf{88.23}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $\textbf{82.18}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.01}}$
RaSa [bai2023rasa] 62.51 83.83 88.09 72.05
+NP $77.04_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.53}}$ $6.66_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.83}} 88$$8.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.76_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.71}}$
+BOP $76.24_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+13.73}}$ $6.74_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.91}} 88$$8.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.29_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.24}}$
MARS [ergasti2024mars] 62.93 84.26 88.65 72.43
+NP $77.26_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.33}}$ $7.30_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.04}} 88$$8.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $82.10_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.67}}$
+BOP $76.57_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+13.64}}$ $7.14_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.88}} 88$$8.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.69_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.26}}$
FMFA [yin2025cross] 59.77 83.22 88.05 70.09
+NP $76.51_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.74}}$ $6.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.36}} 88$$8.05_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.47_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.38}}$
+BOP $75.83_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.06}}$ $6.48_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.26}} 88$$8.05_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.00}}$
+BEP $\mathbf{77.35}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+17.58}}$ $\mathbf{86.59}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.37}}$ $\mathbf{88.05}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $\mathbf{81.95}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.86}}$
1) Naive Prompt (NP), which provides only the textual description and the full-scene image without explicit region information; 2) Box Overlay Prompt (BOP), which visually highlights the candidate region with a red rectangle on the full-scene image; 3) Box Embedded Prompt (BEP), which includes the bounding box coordinates of the candidate region in the multimodal prompt using <box> tags.
As shown in Table V, SceneRanker significantly boosts performance for all methods, with R@1 improvements exceeding above 14.97% with BEP, underscoring its compatibility and effectiveness in refining appearance-based retrieval. Across all methods, BEP consistently outperforms NP and BOP, achieving the highest R@1, suggesting that embedding bounding box coordinates enhances spatial grounding, enabling better alignment with textual descriptions. Notably, BOP even play worse than NP across all baselines, with a representative drop of 1.02% at R@1 and 0.51% at mAP on IRRA with ViT-B/16, indicating that direct visual highlighting may introduce visual distractions that disrupt holistic scene understanding.
Different Multimodal Large Language Models. We evaluate the impact of integrating different 8B-scale MLLMs into the proposed SceneRanker, using the unified naive prompt to mitigate varying capabilities in processing structured inputs. As shown in Table VI, all three models, Qwen-VL 2.5, MiniCPM-V, and InternVL 2.5, consistently improve over the IRRA baseline, with respective R@1 gains of +11.7%, +7.96%, and +14.18%. In addition, we observe a strong correlation between model capacity and performance. Taking InternVL2.5 as an example, scaling from 4B to 8B brings further improvement, with R@1 increasing from 75.21% to 77.55%. In contrast, smaller models below 4B such as InternVL 1B and 2B show significantly lower performance, indicating insufficient multimodal understanding. These results confirm SceneRanker’s effectiveness and generalizability across MLLM architectures with sufficient vision-language reasoning capacity.
@@
TABLE VI: Impact of integrating different MLLMs into SceneRanker. Due to varying MLLM capabilities in processing structured prompts, we adopt Naive Prompt across all models to ensure fair comparison.
@@
IRRA (ViT-L/14) 590 Million 63.37 84.54 88.67 72.71
QwenVL2.5 [bai2025qwen2] 3 Billion 54.57 78.23 88.67 65.04
QwenVL2.5 [bai2025qwen2] 8 Billion 75.07 86.34 88.67 80.45
MiniCPM-V-2_6 [yao2024minicpm] 8 Billion 71.33 84.66 88.67 77.78
InternVL2.5 [chen2024expanding] 1 Billion 3.01 80.14 88.67 26.71
@@
Different Candidate Sizes. In the proposed two-stage framework SA-Person, the candidate set produced by the appearance-based retrieval stage serves as the input for the SceneRanker to perform scene-aware reranking. The size of this set ($K$) governs the trade-off between ensuring the inclusion of the ground-truth match and minimizing the introduction of noisy distractors. To better understand this trade-off and determine an optimal value for ($K$), we conduct a parameter analysis to systematically investigate the impact of different candidate set sizes on the final retrieval performance. As shown in Figure  6, the performance of SceneRanker exhibits a trend of first increasing and then decreasing as the candidate size changes. The best performance is achieved when the candidate size is set to 10. Smaller candidate sets are more likely to exclude the ground-truth instance, while excessively large candidate sets introduce more irrelevant or noisy candidates. Both scenarios negatively impact performance. These results indicate that a moderate candidate pool is optimal, as it balances retaining the ground-truth instance while minimizing noise, ultimately enhancing retrieval accuracy.
VI-E Versatility Across Diverse Retrieval Scenarios
To demonstrate the versatility of SA-Person, we evaluate its performance across three distinct retrieval scenarios: person retrieval in surveillance scenarios using augmented CUHK-SYSU from our preliminary analysis in Section III, person retrieval in varied real-world environments using our SCENEPERSON-13W benchmark, and general image-text retrieval in open-domain contexts using Urban1K [zhang2024long] and a 5K subset of ShareGPT4V [chen2024sharegpt4v].
TABLE VII: SceneRanker Enhancements in Surveillance Scenarios.